OpenStudy (anonymous):

write 100 as a sum of numbers such that the product of those numbers is maximized. @Computer Science

OpenStudy (jamesj):

How many numbers?

OpenStudy (jamesj):

2, 5, 10? An arbitrary finite number?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

I got like \[\sum_{i=1}^{37}2.702702702702....\]

OpenStudy (jamesj):

That's right, because what you're doing is trying to maximize the function \[ f(x) = x^{100/x} \] subject to the constraint that 100/x is an integer. Well if you differentiate that function and set it equal to zero, you'll find that f has a local max at x = e. And as you can see, the x you have in your sum is very close to e.

OpenStudy (anonymous):

so it would be like \[\sum_{i=1}^{37}(2+702\sum_{n=1}^{\infty}{\frac{1}{10^{3n}}})\]

OpenStudy (jamesj):

But why bother with that? All the intuitive content of the term is contained in the simple expression 100/37.

OpenStudy (anonymous):

right